OpenStudy (anonymous):

The length of a rectangle is twice its width. If the area of the retangle is 98 cm^2 (squared), find it's perimeter.

OpenStudy (anonymous):

L = 2W A = L*W P = 2(L+W) You can find the values of the dimensions of the rectangle, and subsequently the perimeter, if you use the above equations :-D

OpenStudy (anonymous):

ok, but what has me confused is the "twice it's width" thing. I don't get these at all.

OpenStudy (anonymous):

It's just saying that if you multiply the width of the rectangle by two, you get the length of the rectangle :)